The following are indirect quotations of the dollar against the Swiss franc, the Japanese yen, and the British pound: $1 = SF1.4814 – 1.4828 $1 = ¥124.26 – 124.37 $1 = £0.6419 – 0.6428 (a) Calculate the midpoints in each rate and the direct quotation of each of the above rates. (b) Calculate the cross rates between (i) the Swiss franc and the Japanese yen, (ii) the Swiss franc and the euro, and (iii) the Japanese yen and the euro.


1
Expert's answer
2021-12-05T18:56:51-0500

a) Swiss franc:

"\frac{1.4814+1.4828}{2}=1.4821"  midpoint; "\frac{1}{1.4821}=0.6747" direct quotation.

Japanese yen:

"\frac{124.26+124.37}{2}=124.32"  midpoint; "\frac{1}{124.32}=0.008" direct quotation.

British pound:

"\frac{0.6419+0.6428}{2}=0.6424"  midpoint; "\frac{1}{0.6424}=1.56" direct quotation.

b) i) "0.6747*124.32=83.88"

ii) "0.6747*0.6424=0.4334"

iii) "0.008*0.6424=0.0051"